OpenStudy (anonymous):

help? Manuel builds and sells wooden crates. He made a table to show how much his crates cost. Which equation represents the situation? Manuel’s Crates Crates 2 4 6 8 Cost 25 45 65 85 A. y – 2 = −10(x – 25) B. y – 4 = 10(x – 45) C. y – 65 = −10(x – 6) D. y – 85 = 10(x – 8)

OpenStudy (anonymous):

@Ishhara9 @StudyGurl14 @sleepyjess

OpenStudy (anonymous):

I'm debating between C and D.

OpenStudy (sleepyjess):

substitute an x value in for x and the corresponding y value for y and see if they equal each other

OpenStudy (anonymous):

I did that in C and D and the answer was always the same on both sides. and there's only one answer that should be correct.

OpenStudy (studygurl14):

Use the slope formula to find the slope. Then use one of the points and plug it into the points-slope formula With point \((x_1,y_1)\) on the line... \(large y-y_1=m(x-x_1)\) m= slope

OpenStudy (studygurl14):

sorry ignore the large part... \(\large y-y_1=m(x-x_1)\)

OpenStudy (sleepyjess):

A. 25 – 2 = −10(2 – 25) B. 25 – 4 = 10(2 – 45) C. 25 – 65 = −10(2 – 6) D. 25 – 85 = 10(2 – 8) A 23=-10(-23) 23=230 nope B 21=10(-43) 21=-430 nope C -40=-10(-4) -40=40 nope D -60=10(-6) -60=-60 yep

OpenStudy (anonymous):

so its D?

OpenStudy (sleepyjess):

yes can you show me your work for C?